OpenStudy (anonymous):

Factor completely: 3x^2 + 10x + 3

OpenStudy (anonymous):

i got (3x + 3)(x + 1)

OpenStudy (anonymous):

@blackprincesssen

OpenStudy (anonymous):

oki hold on

OpenStudy (anonymous):

Factoring: 3x² + 10x + 3 = 0 3x² + 9x + x + 3 = 0 3x(x + 3) + 1(x+3) = 0 (x + 3)(3x + 1) = 0 x + 3 = 0 → x = -3 3x + 1 = 0 → 3x = -1 → x = -⅓

OpenStudy (anonymous):

so its prime?

OpenStudy (anonymous):

@blackprincesssen

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thns

OpenStudy (anonymous):

np